# RefLoRA: Refactored Low-Rank Adaptation for Efficient Fine-Tuning of Large Models

### v1

@@ 3.3 RefLoRA: Refactored low-rank adaptation
Adaptive optimizers such as Adam [25] and AdamW [40] are default to optimizing large models, which adjust the update using the first moment and entrywise second moment estimated from the running average of stochastic gradients. When refactoring $(\mathbf{A}_{t},\mathbf{B}_{t})$ to $(\mathbf{A}_{t}\mathbf{P}_{t},\mathbf{B}_{t}\mathbf{P}_{t}^{-\top})$, the first moment estimator can be transformed accordingly, while the second moment is generally intractable due to the entrywise square. To tackle this challenge, we “refactor back” the low-rank matrices after the GD update in (3) by right-multiplying $\mathbf{P}_{t}^{-1}$ and $\mathbf{P}_{t}^{\top}$. This gives an alternative update
@@
Balanced refactoring $\mathbf{A}_{t}^{\top}\mathbf{A}_{t}=\mathbf{B}_{t}^{\top}\mathbf{B}_{t}$ can be achieved per iteration upon setting $\mathbf{P}_{t}\mathbf{P}_{t}^{\top}=\tilde{\mathbf{S}}_{t}$, as stated in Section 3.2. It is worthwhile pointing out that SVD-based initializations including PiSSA [41] and LoftQ [34] inherently satisfy $\mathbf{A}_{0}^{\top}\mathbf{A}_{0}=\mathbf{B}_{0}^{\top}\mathbf{B}_{0}^{\top}$. When $t$ is small, the balance tends to hold approximately even without refactoring, which partially explains why empirical convergence is fast during the early epochs. Analytically, balanced refactoring maximizes the potential loss reduction as formalized in the ensuing Theorem.
@@
The upper bound 00 stems from the low-rank nature of LoRA, which can be reached when $\nabla\ell(\mathbf{W}_{t})^{\top}\in\operatorname{Null}(\mathbf{B}_{t}^{\top})$ and $\nabla\ell(\mathbf{W}_{t})\in\operatorname{Null}(\mathbf{A}_{t}^{\top})$; i.e., the stationary points of $\mathbf{A}_{t}$ and $\mathbf{B}_{t}$. While this upper bound cannot be improved, RefLoRA minimizes the lower bound, and thus leads to a more effective descent in the loss.
@@
LoRA-Pro [57] $\mathcal{O}(mr^{2}+(m+n+r)r^{2})$ $\mathcal{O}(m^{2}+(m+n+r)r)$
LoRA-RITE [63] $\mathcal{O}((m+n+r)r^{2})$ $\mathcal{O}((m+n+r)r)$
@@
Additional computational overhead induced by RefLoRA is as small as $\mathcal{O}((m+n+r)r^{2})$ in time and $\mathcal{O}(r^{2})$ in memory, thanks to the decoupling of $\nabla\ell(\mathbf{W}_{t})$ in Proposition (2). Compared to the forward/backward overhead $\Omega(mn)$ of LoRA, the extra complexity introduced by RefLoRA is relatively minimal. In contrast, LoRA-Pro [57] suffers from $\mathcal{O}(m^{2}r+(m+n+r)r^{2})$ time and $\mathcal{O}(m^{2}+(m+n+r)r)$ space, whereas LoRA-RITE [63] requires $\mathcal{O}((m+n+r)r)$ memory for polar decomposition. Despite the low complexity of RefLoRA, further curtailing the extra cost can be beneficial for resource-limited applications. The next theorem restricts $\mathbf{S}_{t}$ to a scaled identity $s_{t}\mathbf{I}_{r}$ with $s_{t}>0$, and derives a result analogous to Theorem 3.
@@
Consider (8) confined with $\mathbf{S}_{t}=s_{t}\mathbf{I}_{r},\;s_{t}\in_{++}$. Under Assumptions 1 and 2, it holds
@@
Under Assumptions 1-2 and for $\mathbf{A}_{t}\mathbf{B}_{t}^{\top}=\mathbf{A}_{t}^{\prime}\mathbf{B}_{t}^{\prime\top}$, let $\Delta\tilde{\mathbf{W}}_{t}$ and $\Delta\tilde{\mathbf{W}}_{t}^{\prime}$ be the corresponding weight updates (5) with RefLoRA. It then always holds that $\Delta\tilde{\mathbf{W}}_{t}^{\prime}=\Delta\tilde{\mathbf{W}}_{t}$.
@@
4 Numerical tests
This section evaluates the empirical performance of RefLoRA. All experimental setups including platforms, datasets, models, metrics, and hyperparameters are detailed in Appendix D.
4.1 Matrix factorization
The first numerical test considers low-rank matrix factorization [12]
where $\mathbf{Y}\in^{m\times n}$ is a given low-rank matrix. It can be viewed as applying LoRA to a single-layer model, and training it with whitened data [2; 31]. Figure 2(a) compares the training loss of RefLoRA with LoRA, and another popular approach dubbed ScaledGD [52]—which is tailored particularly for low-rank matrix factorization, and has gained popularity among LoRA variants; see e.g., [65]. Each method is tested with two learning rates $\eta\in\{0.01,0.03\}$. While vanilla LoRA converges slowly with the lower learning rate and diverges with the higher one, RefLoRA remains stable and converges markedly faster under both rates, as corroborated also with observations in Figure 1. Notably, RefLoRA even outperforms ScaledGD, thanks to its balanced update. Figure 2(b) depicts the dynamics of LoRA ($\eta=0.01$) by plotting the Frobenius norms of $\mathbf{A}_{t},\mathbf{B}_{t}$ and their gradients. It is observed that $\|\mathbf{A}_{t}\|_{\mathrm{F}}$ and $\|\mathbf{B}_{t}\|_{\mathrm{F}}$ are highly unbalanced across iterations, and $\|\nabla_{\mathbf{A}_{t}}\ell\|_{\mathrm{F}}$ as well as $\|\nabla_{\mathbf{B}_{t}}\ell\|_{\mathrm{F}}$ exhibit a sharp change when $t$ is small. In comparison, RefLoRA maintains $\tilde{\mathbf{A}}_{t}^{\top}\tilde{\mathbf{A}}_{t}=\tilde{\mathbf{B}}_{t}^{\top}\tilde{\mathbf{B}}_{t},\,\forall t$, which guarantees that $\|\tilde{\mathbf{A}}_{t}\|_{\mathrm{F}}^{2}=\operatorname{tr}(\tilde{\mathbf{A}}_{t}^{\top}\tilde{\mathbf{A}}_{t})=\operatorname{tr}(\tilde{\mathbf{B}}_{t}^{\top}\tilde{\mathbf{B}}_{t})=\|\tilde{\mathbf{B}}_{t}\|_{\mathrm{F}}$. This balance can afford larger learning rates, thus improving the empirical convergence.
@@ 4.2 Natural language understanding
Beyond matrix factorization the evaluation here starts with fine-tuning DeBERTaV3-base [20], a masked language model with 184M parameters, on the General Language Understanding Evaluation (GLUE) benchmark [55]. GLUE contains 8 datasets, providing a general-purpose evaluation for natural language understanding (NLU) [55]. The test setup follows from [22; 66], where LoRA rank is $r=8$, reducing the trainable parameters to 1.33M. We compare RefLoRA and its lightweight variant RefLoRA-S against a suite of LoRA variants, including SOTA methods DoRA [61] and AdaLoRA [66], as well as baselines falling in the same category with RefLoRA, i.e., LoRA-Pro [57] and LoRA-RITE [63]. The results can be found in Table 2. It is worth mentioning that these datasets are relatively small, so that full fine-tuning (FT) is prone to overfitting, thus leading to worse performance compared to PEFT methods. RefLoRA and RefLoRA-S outperform all competitors on 5 out of 8 datasets, and present comparable performance to SOTA approaches on the rest 3 datasets. Overall, RefLoRA achieves the highest average performance, demonstrating more effective optimization via refactoring. In spite of the simplified refactoring, RefLoRA-S maintains competitive performance, i.e., only $0.33\%$ lower than RefLoAR on average, while markedly reducing computational overhead.
@@
4.3 Reasoning
We further extend our numerical experiments to fine-tuning the LLaMA series [53; 54; 16], which are autoregressive language models with 7B and 8B parameters. We tackle commonsense reasoning tasks following the setup in [23; 61]. Training data are aggregated from 8 datasets listed in Table 3, and test sets remain separate for individual evaluation. These reasoning tasks are intended to push the model beyond pattern recognition, requiring commonsense and knowledge to make proper inferences. The baselines are chosen as DoRA [61], PrecLoRA [65], and NoRA+ [31]. Note that the latter two approaches are variants of ScaledGD [52], sharing similar complexity with RefLoRA. The accuracy comparison is summarized in Table 3. Both RefLoRA and RefLoRA-S consistently outperform other PEFT methods in 5 out of 6 settings. Even under lower-rank configurations $r=16$, both RefLoRA and RefLoRA-S continue to lead or match top-performing approaches, underscoring their parameter efficiency and robustness. These results demonstrate the effectiveness of the proposed RefLoRA(-S), and underscore the potential of optimal refactoring.
@@ 4.4 Convergence and complexity comparison
Lastly, we evaluate the convergence behavior and computational efficiency of RefLoRA in comparison with LoRA [22], LoRA-Pro [57], and LoRA-RITE [63]. These tests are conducted on the MRPC subset of the GLUE benchmark using DeBERTaV3-base as the backbone model. For fairness, the learning rate is set to $\eta=4\times 10^{-4}$ across all methods. Figure 3(a) depicts the loss $\ell(\mathbf{W}_{t})$ over $10$ fine-tuning epochs. The loss of RefLoRA declines more rapidly and exhibits less fluctuation than the other three methods, which ultimately achieves the lowest value approaching 00. This confirms stability and convergence speed, on par with our theoretical insights in Section 3. The sharper descent in the loss suggests improved optimization trajectories enabled by RefLoRA’s principled refactoring.
The comparison of computational overhead is presented in Figure 3(b). Time complexity is reflected via the fine-tuning throughput (iterations per second; higher is better), while space complexity is measured in GPU memory occupation (lower is better). For better visualization, the vertical axes start with a non-zero value. The plot reveals that RefLoRA and RefLoRA-S respectively showcase $88.5\%$ and $98.7\%$ throughput compared to LoRA, at the additional memory cost of $132$MB and $<1$MB. In contrast, the throughput of LoRA-Pro and LoRA-RITE are $60.2\%$ and $72.6\%$ of LoRA, requiring $134$MB and $140$MB extra memory. This is consistent with our complexity analysis in Table 1.
@@ Appendix A Additional related work
LoRA variants. LoRA has been extended in several directions. For example, recent works further reduce LoRA’s trainable parameters [27, 37, 14, 18, 26]. As these methods can achieve performance comparable to that of LoRA, they indirectly suggest that the expressiveness of LoRA has not been fully exploited. Another line of work [10, 34] incorporates quantization in LoRA to reduce memory footprint and computational overhead. There are also works that broaden the applicability of LoRA to pre-training LLMs by e.g., sequentially chaining LoRA modules [35, 39]. As our method enhances the efficiency of LoRA training, we expect that it can be seamlessly integrated to such settings as well. Going beyond LoRA, there are also approaches for fine-tuning LLMs in a parameter efficient manner such as [33, 28, 38, 60]. Moreover, [30] shows that applying sharpness-aware minimization (SAM) [13, 29] on LoRA promotes balance between $\mathbf{A}_{t}$ and $\mathbf{B}_{t}$. These approaches are orthogonal to our work.
@@
Future directions. Due to limited computational resources, our evaluation currently deals with models having reasonably large scale, e.g., LLaMA3-8B. Our future work will include scaling RefLoRA to even larger models, such as those with 30B parameters. Another promising direction is to integrate RefLoRA with sequentially chaining, namely  [35]. This will further broaden the applicability of RefLoRA for pre-training LLMs.

### v2

@@ 3.3 RefLoRA: Refactored low-rank adaptation
Adaptive optimizers such as Adam Adam and AdamW AdamW are default to optimizing large models, which adjust the update using the first moment and entrywise second moment estimated from the running average of stochastic gradients. When refactoring $(\mathbf{A}_{t},\mathbf{B}_{t})$ to $(\mathbf{A}_{t}\mathbf{P}_{t},\mathbf{B}_{t}\mathbf{P}_{t}^{-\top})$, the first moment estimator can be transformed accordingly, while the second moment is generally intractable due to the entrywise square. To tackle this challenge, we “refactor back” the low-rank matrices after the GD update in (3) by right-multiplying $\mathbf{P}_{t}^{-1}$ and $\mathbf{P}_{t}^{\top}$. This gives an alternative update
@@
Balanced refactoring $\mathbf{A}_{t}^{\top}\mathbf{A}_{t}=\mathbf{B}_{t}^{\top}\mathbf{B}_{t}$ can be achieved per iteration upon setting $\mathbf{P}_{t}\mathbf{P}_{t}^{\top}=\tilde{\mathbf{S}}_{t}$, as stated in Section 3.2. It is worthwhile pointing out that SVD-based initializations including PiSSA PiSSA and LoftQ LoftQ inherently satisfy $\mathbf{A}_{0}^{\top}\mathbf{A}_{0}=\mathbf{B}_{0}^{\top}\mathbf{B}_{0}^{\top}$. When $t$ is small, the balance tends to hold approximately even without refactoring, which partially explains why empirical convergence is fast during the early epochs. Analytically, balanced refactoring maximizes the potential loss reduction as formalized in the ensuing Theorem.
@@
The upper bound 0 stems from the low-rank nature of LoRA, which can be reached when $\nabla\ell(\mathbf{W}_{t})^{\top}\in\operatorname{Null}(\mathbf{B}_{t}^{\top})$ and $\nabla\ell(\mathbf{W}_{t})\in\operatorname{Null}(\mathbf{A}_{t}^{\top})$; i.e., the stationary points of $\mathbf{A}_{t}$ and $\mathbf{B}_{t}$. While this upper bound cannot be improved, RefLoRA minimizes the lower bound, and thus leads to a more effective descent in the loss.
@@
LoRA-Pro LoRA-Pro $\mathcal{O}(m^{2}r+(m+n+r)r^{2})$ $\mathcal{O}(m^{2}+(m+n+r)r)$
LoRA-RITE LoRA-RITE $\mathcal{O}((m+n+r)r^{2})$ $\mathcal{O}((m+n+r)r)$
@@
Additional computational overhead induced by RefLoRA is as small as $\mathcal{O}((m+n+r)r^{2})$ in time and $\mathcal{O}(r^{2})$ in memory, thanks to the decoupling of $\nabla\ell(\mathbf{W}_{t})$ in Proposition (2). Compared to the forward/backward overhead $\Omega(mn)$ of LoRA, the extra complexity introduced by RefLoRA is relatively minimal. In contrast, LoRA-Pro LoRA-Pro suffers from $\mathcal{O}(m^{2}r+(m+n+r)r^{2})$ time and $\mathcal{O}(m^{2}+(m+n+r)r)$ space, whereas LoRA-RITE LoRA-RITE requires $\mathcal{O}((m+n+r)r)$ memory for polar decomposition. Despite the low complexity of RefLoRA, further curtailing the extra cost can be beneficial for resource-limited applications. The next theorem restricts $\mathbf{S}_{t}$ to a scaled identity $s_{t}\mathbf{I}_{r}$ with $s_{t}>0$, and derives a result analogous to Theorem 3.
@@
Consider (8) confined with $\mathbf{S}_{t}=s_{t}\mathbf{I}_{r},\;s_{t}\in\mathbb{R}_{++}$. Under Assumptions 1 and 2, it holds
@@
Under Assumptions 1-2 and for any $\mathbf{A}_{t}^{\prime}\mathbf{B}_{t}^{\prime\top}=\mathbf{A}_{t}\mathbf{B}_{t}^{\top}$, let $\Delta\tilde{\mathbf{W}}_{t}^{\prime}$ and $\Delta\tilde{\mathbf{W}}_{t}$ be the corresponding weight updates (5) with RefLoRA. It then always holds that $\Delta\tilde{\mathbf{W}}_{t}^{\prime}=\Delta\tilde{\mathbf{W}}_{t}$.
@@
4 Numerical tests
This section evaluates the empirical performance of RefLoRA. All experimental setups including platforms, datasets, models, metrics, and hyperparameters are detailed in Appendix D. Our implementation is available at https://github.com/zhangyilang/RefLoRA.
4.1 Matrix factorization
The first numerical test considers low-rank matrix factorization MF-GD
where $\mathbf{Y}\in\mathbb{R}^{m\times n}$ is a given low-rank matrix. It can be viewed as applying LoRA to a single-layer model, and training it with whitened data arora2018 ; NoRA . Figure 2(a) compares the training loss of RefLoRA with LoRA, and another popular approach dubbed ScaledGD ScaledGD —which is tailored particularly for low-rank matrix factorization, and has gained popularity among LoRA variants; see e.g., PrecLoRA . Each method is tested with two learning rates $\eta\in\{0.01,0.03\}$. While vanilla LoRA converges slowly with the lower learning rate and diverges with the higher one, RefLoRA remains stable and converges markedly faster under both rates, as corroborated also with observations in Figure 1. Notably, RefLoRA even outperforms ScaledGD, thanks to its balanced update. Figure 2(b) depicts the dynamics of LoRA ($\eta=0.01$) by plotting the Frobenius norms of $\mathbf{A}_{t},\mathbf{B}_{t}$ and their gradients. It is observed that $\|\mathbf{A}_{t}\|_{\mathrm{F}}$ and $\|\mathbf{B}_{t}\|_{\mathrm{F}}$ are highly unbalanced across iterations, and $\|\nabla_{\mathbf{A}_{t}}\ell\|_{\mathrm{F}}$ as well as $\|\nabla_{\mathbf{B}_{t}}\ell\|_{\mathrm{F}}$ exhibit a sharp change when $t$ is small. In comparison, RefLoRA maintains $\tilde{\mathbf{A}}_{t}^{\top}\tilde{\mathbf{A}}_{t}=\tilde{\mathbf{B}}_{t}^{\top}\tilde{\mathbf{B}}_{t},\,\forall t$, which guarantees that $\|\tilde{\mathbf{A}}_{t}\|_{\mathrm{F}}^{2}=\operatorname{tr}(\tilde{\mathbf{A}}_{t}^{\top}\tilde{\mathbf{A}}_{t})=\operatorname{tr}(\tilde{\mathbf{B}}_{t}^{\top}\tilde{\mathbf{B}}_{t})=\|\tilde{\mathbf{B}}_{t}\|_{\mathrm{F}}$. This balance can afford larger learning rates, thus improving the empirical convergence.
@@ 4.2 Natural language understanding
Beyond matrix factorization the evaluation here starts with fine-tuning DeBERTaV3-base DeBERTaV3 , a masked language model with 184M parameters, on the General Language Understanding Evaluation (GLUE) benchmark wang2018glue . GLUE contains 8 datasets, providing a general-purpose evaluation for natural language understanding (NLU) wang2018glue . The test setup follows from LoRA ; AdaLoRA , where LoRA rank is $r=8$, reducing the trainable parameters to 1.33M. We compare RefLoRA and its lightweight variant RefLoRA-S against a suite of LoRA variants, including SOTA methods DoRA DoRA and AdaLoRA AdaLoRA , as well as baselines falling in the same category with RefLoRA, i.e., LoRA-Pro LoRA-Pro and LoRA-RITE LoRA-RITE . The results can be found in Table 2. It is worth mentioning that these datasets are relatively small, so that full fine-tuning (FT) is prone to overfitting, thus leading to worse performance compared to PEFT methods. RefLoRA and RefLoRA-S outperform all competitors on 5 out of 8 datasets, and present comparable performance to SOTA approaches on the rest 3 datasets. Overall, RefLoRA achieves the highest average performance, demonstrating more effective optimization via refactoring. In spite of the simplified refactoring, RefLoRA-S maintains competitive performance, i.e., only $0.33\%$ lower than RefLoAR on average, while markedly reducing computational overhead.
@@
4.3 Commonsensing reasoning
We further extend our numerical experiments to fine-tuning the LLaMA series llama ; llama2 ; Llama3 , which are autoregressive language models with 7B and 8B parameters. We tackle commonsense reasoning tasks following the setup in hu2023llm ; DoRA . Training data are aggregated from 8 datasets listed in Table 3, and test sets remain separate for individual evaluation. These reasoning tasks are intended to push the model beyond pattern recognition, requiring commonsense and knowledge to make proper inferences. The baselines are chosen as DoRA DoRA , LoRA-RITE LoRA-RITE , PrecLoRA PrecLoRA , and NoRA+ NoRA . Note that the latter two approaches are variants of ScaledGD ScaledGD , sharing similar complexity with RefLoRA. The accuracy comparison is summarized in Table 3. Both RefLoRA and RefLoRA-S consistently outperform other PEFT methods in 5 out of 6 settings. Even under lower-rank configurations $r=16$, both RefLoRA and RefLoRA-S continue to lead or match top-performing approaches, underscoring their parameter efficiency and robustness. These results demonstrate the effectiveness of the proposed RefLoRA(-S), and underscore the potential of optimal refactoring.
Table 4: Fine-tuning loss ($\downarrow$) for subject-driven image generation on DreamBooth.
Loss LoRA LoRA-Pro LoRA-RITE RefLoRA
[…]
Figure 3: Images generated from Stable Diffusion fine-tuned with different approaches.
4.4 Subject-driven image generation with diffusion models
Akin to LoRA, RefLoRA can be seamlessly integrated into a wide range of larger models beyond LLMs. Further numerical tests are conducted on a subject-driven image generation task DreamBooth with Stable Diffusion v1.4 stable-diffusion . The goal is to fine-tune a diffusion model using a few user-provided images so that it can generate the same object in various contexts. Specifically, the model is fine-tuned on a set of images labeled “a photo of sks dog,” and subsequently evaluated by generating images under the prompt “a sks dog eating nachos.” LoRA adapters with rank $r=4$ are attached to the U-Net component of Stable Diffusion, and experimental setups including hyperparameters are default to those in DreamBooth . The fine-tuning losses for LoRA, LoRA-Pro, LoRA-RITE, and RefLoRA are summarized in the table 4, where average and standard deviation are calculated over 3 random runs. It is seen that RefLoRA achieves 14.0% , 13.1%, and 9.5% improvements over LoRA and LoRA-Pro, and LoRA-RITE, respectively. In addition to quantitative gains in loss, we also observe noticeable improvements in image quality; see Figure 3. The generations from RefLoRA-tuned models show markedly clearer details and better object fidelity, particularly in the mouth and tongue regions, where features are often distorted in outputs from other three baselines.
@@ 4.5 Convergence and complexity comparison
Lastly, we evaluate the convergence behavior and computational efficiency of RefLoRA in comparison with LoRA LoRA , LoRA-Pro LoRA-Pro , and LoRA-RITE LoRA-RITE . These tests are conducted on the MRPC subset of the GLUE benchmark using DeBERTaV3-base as the backbone model. For fairness, the learning rate is set to $\eta=4\times 10^{-4}$ across all methods. Figure 4(a) depicts the loss $\ell(\mathbf{W}_{t})$ over $10$ fine-tuning epochs. The loss of RefLoRA(-S) declines more rapidly and exhibits less fluctuation than the other three methods, which ultimately achieves the lowest value approaching 0. This confirms stability and convergence speed, on par with our theoretical insights in Section 3. The sharper descent in the loss suggests improved optimization trajectories enabled by RefLoRA’s principled refactoring.
The comparison of computational overhead is presented in Figure 4(b). Time complexity is reflected via the fine-tuning throughput (iterations per second; higher is better), while space complexity is measured in GPU memory occupation (lower is better). For better visualization, the vertical axes start with a non-zero value. The plot reveals that RefLoRA and RefLoRA-S respectively showcase $88.5\%$ and $98.7\%$ throughput compared to LoRA, at the additional memory cost of 132132MB and <1<1MB. In contrast, the throughput of LoRA-Pro and LoRA-RITE are $60.2\%$ and $72.6\%$ of LoRA, requiring 134134MB and 140140MB extra memory. This is consistent with our complexity analysis in Table 1. Extended comparisons scaling up to 27B models are offered in Appendix E.
@@ 5 Conclusion and outlook
Acknowledgments
This work is partly supported by NSF grants 2220292, 2212318, 2312547, and 2332547. B. Li is supported by Swiss National Science Foundation (SNSF) Project Funding No. 200021-207343.
@@ Appendix A Additional related work
LoRA variants. LoRA has been extended in several directions. For example, recent works further reduce LoRA’s trainable parameters kopiczko2023vera ; lingam2024svft ; gao2024 ; hao2024flora ; koike2025quantum . As these methods can achieve performance comparable to that of LoRA, they indirectly suggest that the expressiveness of LoRA has not been fully exploited. Another line of work QLoRA ; LoftQ incorporates quantization in LoRA to reduce memory footprint and computational overhead. There are also works that broaden the applicability of LoRA to pre-training LLMs by e.g., sequentially chaining LoRA modules lialin2023relora ; liu2025cola . As our method enhances the efficiency of LoRA training, we expect that it can be seamlessly integrated to such settings as well. Going beyond LoRA, there are also approaches for fine-tuning LLMs in a parameter efficient manner such as prefix-tuning ; prompt-tuning ; liu2023parameter ; wu2024reft . Moreover, li2024bar shows that applying sharpness-aware minimization on LoRA promotes balance between $\mathbf{A}_{t}$ and $\mathbf{B}_{t}$. These approaches are orthogonal to our work.
@@
Future directions. Due to limited computational resources, our evaluation currently deals with models having reasonably large scale, e.g., LLaMA3-8B. Our future work will include scaling RefLoRA to even larger models, such as those with 30B parameters. Another promising direction is to integrate RefLoRA with sequentially chaining, namely  lialin2023relora . This will further broaden the applicability of RefLoRA for pre-training LLMs.